Accelerator Update for June 6, 2008 - June 9, 2008
Friday June 6
The day shift began with Operations monitoring Tevatron (TeV) store 6200 (188.0E30) and the Recycler stash (224.7E10), with the Antiproton Source (Pbar) stacking (12.5E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 11:16 AM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 & 8 to also trip off. Everything reset.
At 1:08 PM, a Linac RF station (LRF3) tripped off. It reset.
Two Pbar experts began momentum aperture studies at 2:36 PM.
Operations reported at 3:57 PM that the Recycler had stashed antiprotons five times during the day shift.
Permission (ORC) to send beam to the new MTest was granted around 4 PM.
A MI expert began a study period at 4:29 PM.
Operations reported at 11:53 PM that the Recycler had stashed two times during the evening shift.
Saturday June 7
The midnight shift began with Operations monitoring store 6200 (57.86E30) and the stash (405.1E10), with Pbar stacking (112.1E10), and with BNB taking beam.
Operations turned off the Pbar beam at midnight in preparation for work in the AP0 water cage.
At 1:15 AM, a Booster RF station (BRF10) tripped off. It reset.
Operations terminated store 6200 at 6:03 AM.
Operations prepared D0 and CDF for access by 6:27 AM.
FESS personnel began work on the TeV sector A1 pond water strainer at 7:09 AM.
FESS personnel reported at 8:04 AM that they had completed their work on the sector F4 pond water pump. They continued working on the other pond water pumps around the TeV all day.
Operations reported at 8:13 AM that while operators and a TeV expert were narrowing down the TeV ground fault location, the ground fault went away.
Operations prepared the TeV for an access at 8:17 AM.
At 8:32 AM, the water group reported finding the TeV sector F4 heat exchanger clogged with mussels. All the other heat exchangers around the TeV ring will also require cleaning at some future date.
At 10:32 AM, Operations reported a problem with the antiproton stacktail. Operators investigated and found it was due to insufficient stacktail power.
Operators reported at 10:57 AM that they had walked through the transfer hall and found ice balls at TeV sector A21. This could have been the cause of the TeV ground fault, but that could not be confirmed.
At 11:47 AM, a Linac Klystron RF station (KRF5) tripped off. It reset.
D0 and the TeV came out of access around 1:06 PM.
Operations put the TeV through a dry squeeze at 2:52 PM.
Operations put the TeV through a wet squeeze at 3:24 PM.
At 3:28 PM, operators responded to a TeV sector C1 air compressor trip. They reported that the blowdown valve was stuck open.
At 3:29 PM, Operations reported a problem with sending beam to MTest; a septa (FSEPTV) was off due to a vacuum problem. A SY expert said that the beam could be sent to MTest temporarily by special tuning.
At 3:58 PM, Operations reported that the Booster low-level RF room temperature had alarmed. Operations requested FESS to install a portable AC unit.
Operations put the TeV into shot setup at 4:58 PM.
At 5:24 PM, one of the two AC units at SciBooNE tripped off. An AC tech was called in to repair the unit.
At 5:41 PM, MIRF13 tripped off and would not reset. Operators bypassed the station and called in a RF expert.
At 6:11 PM, Operations turned off the stacking due to a large stack (186E10) that would not tolerate the quicker rep-rate.
At 6:20 PM, Operations reported a glitch that tripped off the Booster West Anode Power Supply (WAPS) and various Pbar devices. Everything reset and no beam was lost.
At 6:36 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to also trip. Everything reset.
At 6:50 PM, Operations reported that beam was falling out of the Accumulator rapidly. Operations contacted Pbar experts.
At 6:58 PM, Operations reported that the Pbar dry engine tripped off. Cryo system experts at CHL reported that their lights flickered at the same time as the dry engine tripped off, but operators could find no indication of a power glitch.
At 7:05 PM, a Pbar expert began investigating the loss of antiprotons, but then decided the best thing to do was to transfer the remaining beam to the Recycler.
Operations established store 6203 at 7:53 PM with an initial luminosity of 288.8E30.
Cryo system experts had the Pbar dry engine up and running by 8:13 PM. They reported that it looked like the engine had tripped on a power glitch.
Operations established beam to the MTest experiment at 8:01 PM.
At 9:04 PM, KRF1 tripped off. It reset.
At 9:11 PM, the P1 transfer line beam permit tripped off due to the autotune program setting a Power supply (HT701) to a bad value. This was the first of three trips in about an hour. An operator adjusted the flattop value for a Lambertson (I:LAM52) twice and then HT701 went to its normal setting.
At 11:18 PM, BRF15 tripped off. It reset.
Operations disabled the MTest beam at 11:45 PM.
Sunday June 8
The midnight shift began with Operations monitoring store 6203 and the stash, with Pbar stacking, and with BNB taking beam.
At 6:15 AM, LRF3 tripped off. It reset.
Operations reported at 7:59 AM that the Recycler had stashed eight times during the last 12 hours.
At 8:55 AM, KRF7 tripped off. It reset.
Operations terminated store 6203 at 9:03 AM.
Operations put the TeV into shot setup at 9:15 AM.
A MI expert began a coalescing check at 9:43 AM.
At 10:44 AM, BRF15 tripped off. It reset.
Operations established store 6204 at 11:02 AM with an initial luminosity of 279.6E30.
Operations established beam to MTest at 11:55 AM.
At 12:48 PM, Operations reported that a TeV sector A0 compressor was drawing more current than normal. Operators and a cryo system expert investigated. The Cryo Coordinator told CHL to shut the compressor down before it caused a yard breaker trip. Experts will come in and investigate the problem.
At 3:53 PM, LRF1 tripped off and would not reset. Operations called in Linac experts.
At 4:16 PM, BRF19 tripped off. A Booster vacuum pump (IPRF17) also tripped off and caused BRF7 & 8 to trip off. Everything eventually reset.
At 4:30 PM, store 6204 aborted; there was no quench. The TeV sector C3 vacuum pulled the abort. After some investigation, Operations reported that the same thing had happened last year at C3 and the problem was with the C200 module.
Operations loaded an energy conservation timeline into the Time Line Generator (TLG) at 5:02 PM.
At 6:14 PM, Operations reported that MI kickers (KPS1A, B, and C) kept tripping off. They all reset, but then tripped off again within 15 minutes. Operations contacted MI experts.
Operations reported at 6:39 PM that the Linac experts had LRF1 back online. The experts said they had found and replaced a blown 12-amp fuse.
Operations disabled the MTest beam at 6:44 PM.
Operators reported at 6:50 PM that they had replaced the 200 module for TeV crate C3.
Operations put the TeV through a dry squeeze at 6:52 PM.
At 7:05 PM, Operations reported that the MI injection kickers continued to trip off. Operators visited MI-10, but everything seemed fine.
At 7:47 PM, the MI beam permit tripped off on a vacuum beam valve (I:BV846) closing. Everything reset.
Operations reported at 8:02 PM that the Recycler had stashed three times during the last 12 hours.
At 9:13 PM, a Recycler RF station (AR5KW4) tripped off. It reset.
Operations put the TeV through a wet squeeze at 9:19 PM.
Operations put the TeV into shot setup at 9:33 PM.
At 10:09 PM, a MI kicker (I:KPS5S) tripped off. It reset.
Operations established store 6207 at 11:10 PM with an initial luminosity of 186.0E30.
Monday June 9
The midnight shift began with Operations monitoring store 6207 and the stash, with Pbar stacking, and with BNB taking beam.
At 12:11 AM, store 6207 aborted. There was no quench. The initial cause indicated bad beam positions at D0, but this was really due to a lightning strike.
At 1:14 AM, MIRF12 tripped off. It reset.
At 2:05 AM, a Recycler RF station (AR5KW4) tripped off for the first of two times in three hours. It reset.
Operations put the TeV through a wet squeeze at 4:03 AM.
Operations put the TeV into shot setup at 4:25 AM.
Operations established store 6210 at 5:44 AM with an initial luminosity of 121.6E30.
Operations reported at 8 AM that the Recycler had stashed six times during the last 12 hours.
The Plans for Monday
The plans for today and the are to stack, stash, and store. The Run Coordinator reported that tomorrow, 6/10/08, there might be a two hour access into MI to get some NuMI equipment out of the tunnel. He also might allow other jobs that could be completed in two hours or less.
Comments and Suggestions
|last modified 06/09/2008 email Fermilab|
|Security, Privacy, Legal|